OpenStudy (anonymous):

Find the area of a square if the perimeter is equal to 40 in.???

OpenStudy (unklerhaukus):

Area of a square = L^2 Perimeter of a square = 4L

OpenStudy (anonymous):

let x for sides P=4x 40=4x x=40/4 x=10

OpenStudy (anonymous):

A=10*10 A=100

OpenStudy (anonymous):

100 SQUARE INCHESASASASASAS...

OpenStudy (anonymous):

\[A=100^{2}\]in